OpenStudy (anonymous):

l4x-3l=9 solve and graph

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

this is a system:{ 4x-3=9 {3-4x=9 since the right sides are = we can do this: 4x-3=3-4x 8x=6 x=6/8 i guess this is the solution :) and i can't graph it here O.o

OpenStudy (anonymous):

hi angela , your approach leaves out a major condition of the mod function:when its argument is greater than 0 only a single equation exists. therefor both equations can never simultaneously exist. for soltero, Hence the correct answer is that there are 2 solutions 1 corresponding to each region of the function ( i.e. >0 & <0). solving the equations separately we get x = 3,x = -6/4 . As for the graph it looks like a 'V' with its vertex at 3/4. I Hope my answer was helpful.

OpenStudy (angela210793):

Oops...i don't get why i have solved tht in that way when this has only 1 variable...:$:$ It was mornig,i hadn't had breakfast....maybe this is the reason :) Thanks :)